OpenStudy (dumbsearch2):

find the discriminant x^2 = 4

hartnn (hartnn):

first write it in the form of \(ax^2+bx+c=0\) find, a,b,c then discriminant is just \(\Large D =b^2-4ac\)

OpenStudy (nincompoop):

your equation in full form looks like this \[1x^2+0x-4=0\] \[a=1, b=0, c=-4\]

OpenStudy (nincompoop):

now plug and chug

hartnn (hartnn):

we need not solve it, only find the discriminant...and no, thats still not ok....

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

@chathuranga , be care ful! -2 is another answer too as (-2)^2 = 4 too